OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (anonymous):

Equation: y-y1=m(x-x1) Knowns: m=2 (x1,y1)=(3, -2) Attempt at solving: This is your part. Plug the values given to the equation that I wrote.

OpenStudy (anonymous):

\[at equation(Y-Y1)=m(X-X1); then:Y-(-2)=2(X-3) ; then:Y+2=2X-6\] the equation is "Y-2X=-8"

OpenStudy (anonymous):

I got 2 answers and neither are a choice..

OpenStudy (anonymous):

Oh! i got it now! I added 3 instead of subtract it.